Rohan have some books and some tables .If he puts 9 books on each table he will have two tables left vacant.If he put 6 books on each table he will have 3 books left over .Find the numbers of books

Answer by Edwin McCravy(20054)

Suppose there are B books and T tables

If he puts 9 books on each table he will have two tables left vacant.
B = 9(T-2)  

If he put 6 books on each table he will have 3 books left over .Find the numbers of books
6T+3 = B

6T+3 = 9(T-2)
6T+3 = 9T-18
  21 = 3T
   7 = T

6T+3 = 6(7)+3 = 42+3 = 45 = B

So there are 7 tables and 45 books.

Edwin